OpenStudy (anonymous):

If given f(x) = x + 4 and g(x) = –2x2 + 3x – 1, what is (f – g)(x)?

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

Hey hey :)

OpenStudy (anonymous):

hi hi :D

OpenStudy (anonymous):

thought you were busy tbh

OpenStudy (anonymous):

Always available to help! So this notation can be confusing, it actually just means \[(f-g)(x) \implies f(x)-g(x)\]

OpenStudy (anonymous):

\[(f-g)(x) = x+4-(-2x^2+3x-1)\]

OpenStudy (anonymous):

As always you can simplify and what not

OpenStudy (anonymous):

alright thank you again :)

OpenStudy (anonymous):

Np ^.^

OpenStudy (anonymous):

so would the answer be (f – g)(x) = –2x2 – 2x + 5 ?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Close, you should have positive 2x^2

OpenStudy (anonymous):

ah so (f – g)(x) = 2x^2 – 2x + 5

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Yes, that looks good :)

OpenStudy (anonymous):

alright thanks again

OpenStudy (anonymous):

Np